Are the statement true or false?give reason for your answers: the function,f(x) =sin^2x is uniformly continuous in the interval [0,π]

1
Expert's answer
2022-01-04T11:20:45-0500

Since the function "f(x)=\sin^2x" is a composition of the elementary continuous function "g(x)=x^2" and "h(x)=\sin x," we conclude that the function "f" is continuous in the interval "[0,\u03c0]."

The Heine-Cantor theorem asserts that every continuous function on a compact set is uniformly continuous. In particular, if a function is continuous on a closed bounded interval of the real line, it is uniformly continuous on that interval. It follows that the function "f(x)=\sin^2x" is uniformly continuous in the interval "[0,\u03c0]."


Answer: true